OpenStudy (timaashorty):

what values for Ɵ ( 0 ≤ Ɵ ≤ 2 π) satisfy the equation? Tan^2 = - 3/2 sec Ɵ

OpenStudy (anonymous):

Check your typing on the equation?

OpenStudy (timaashorty):

oops > Tan^2 ÆŸ = - 3/2 sec ÆŸ

OpenStudy (timaashorty):

Nope, that's a diff. equation..

OpenStudy (anonymous):

\[\tan^{2}(\theta) = \frac{-3}{2}\sec(\theta) ?\]

OpenStudy (timaashorty):

yes [:

OpenStudy (timaashorty):

@satellite73

OpenStudy (tkhunny):

It's a Quadratic Problem, Folks. Convert tangent to secant and you should see it.

OpenStudy (timaashorty):

can you please explain further?

OpenStudy (anonymous):

what trig identity do you know?

OpenStudy (tkhunny):

Maybe \(\tan^{2}(x) + 1 = \sec^{2}(x)\)? There's a ton of these. You must work on memorizing more and more of them, They will serve you well.